Archive for March, 2010

North Paddington Exchange

Posted: Wednesday, March 31st, 2010 at 17:20 by Ryan Berrisford

We are currently aware of an issue in the North Paddington Exchange affecting leased lines, DSL and PSTN based circuits. Damage to the exchange hasĀ  reportedly been caused by flooding and fire although the sequence of the events is not clear at present. We have been advised though that the fire is now out and flooding has been stopped. Water pumps are currently on site to clear water from the exchange, at which point engineers will be able to fully assess the damage.

An initial ETR has been set as 12:00 on the 02/04/2010 although further updates will follow as and when they become available.

  1. Marc Woolfenden says:

    We have been advised by our suppliers that power engineers have been working on site over the course of the night. At this time power has been restored to approximately 75 percent of devices in the exchange.

    Upon full restoration of power to the site Openreach engineers will begin to re-establish priority services on the network platform.

    Further updates will be provided as more information becomes availiable.

  2. Marc Woolfenden says:

    Our suppliers have advised that work on this issue is still ongoing. We have however been informed that service has been restored to a portion of the leased line and DSL circuits which were impacted by this incident.

    Further updates will be provided as more information becomes availiable.

  3. Ryan Berrisford says:

    Our suppliers have advised that work on this issue is still ongoing, however power and the majority of services have now been restored.

    Engineers are working on the small portion of services still affected and further updates will follow as and when we receive them.

  4. Ryan Berrisford says:

    This issue has now been confirmed resolved.

    Any users that remain affected after retesting the service are advised to report this via the normal support channels.

Midlands based 21CN DSL connections

Posted: Wednesday, March 31st, 2010 at 09:22 by Richard Partridge

We are currently aware of a Major Service Outage on BT’s network which is affecting a number of 21CN based DSL connections in the Midlands region. This outage is the result of a failed BT Planned Engineering Works event which was in progress during the early hours of this morning.

BT are currently working to resolve this issue, however no ETR has been announced as of yet.

Further updates will be posted as the information is received from BT.

  1. Marc Woolfenden says:

    A large number of sessions have been restored over the course of the morning, however a small amount of users remain affected.

    We have been advised that work is still ongoing however there is no ETR as some resources have been diverted to an emergency issue at London Paddington following a fire in that exchange which has impacted priority services.

    Further updates will follow as more information becomes availiable.

  2. Marc Woolfenden says:

    This issue has now been confirmed resolved.

    Users who remain affected should power down their ADSL equipment for a period of 5 minutes in order to make a clean reconnection attempt. Any issues persisting beyond this should be reported to the Support Desk via the normal channels.

Leeds to Manchester

Posted: Friday, March 26th, 2010 at 09:05 by Steve Lalonde

Our link between Leeds and Manchester is currently down. We are investigating the cause with our supplier. Traffic is being rerouted via alternate paths, some increase in latency may be seen.

  1. Phil Davies says:

    Our Supplier has informed us that a fiber break has been located in the Manchester area and engineers are on site doing the prep work to repair the break. No repair time has been given as yet but any updates will be posted as soon as we recieve them.

  2. Phil Davies says:

    The circuit is now restored between Manchester and Leeds however it is still considered at risk untill all work has been completed and the fault has been closed off.

WBC user authentication

Posted: Monday, March 15th, 2010 at 19:26 by Steve Lalonde

We have been experiencing authentication problems for some WBC users over the last 2 hours.

This issue has now been resolved and all affected users have reconnected.

  1. Marc Woolfenden says:

    There is now an outage report available for this issue. It can be found here.

Outage: Wolverhampton to Birmingham

Posted: Monday, March 15th, 2010 at 15:34 by Martyn Haynes

A link between Wolverhampton and Birmingham has gone down. We are investigating with our supplier. Traffic is being rerouted via alternate paths, some increase in latency may be seen.

  1. Martyn Haynes says:

    We have been advised that vandalism has been reported by Network Rail on a bridge which carries these fibres, engineer has been despatched to find breakage and is due on-site within the hour.

    Further updates will be made as and when they are available.

  2. Martyn Haynes says:

    We have been informed that engineers should now be on site, with further updates to follow.

  3. Steve Lalonde says:

    Service was restored at 21:20 traffic is now back on the preferred paths.

Emergency maintenance: VOIP platform

Posted: Monday, March 15th, 2010 at 11:35 by David Derrick

Between 06:00 and 06:30 tomorrow (16 March 2010) there will be a 15 minute downtime for the Entanet VOIP system. This is to enable the replacement of a faulty UPS. The system will be unavailable during that time.

  1. David Derrick says:

    This work has been completed successfully.

Outage: Power issues at Telford HQ

Posted: Saturday, March 13th, 2010 at 12:17 by Neil Watson

We are aware that we have experienced power issues this morning at out Telford HQ that has affected hosting in our Telford based data centres as well as any connectivity that was routed through Telford. The issues have also affected our VoIP service, hence customers will have experienced problems with inbound and outbound calls, which also affected our support service as no calls could be received.

Power has been restored to the site and engineers are present and working to restore services. Initial indications are that the power failed following a scheduled generator test, however it has yet to be confirmed whether this was the cause of the outage. Will will update as more information becomes available.

  1. Marc Woolfenden says:

    There is now an outage report available for this issue. It can be found here.

Outage: Wolverhampton to Birmingham

Posted: Friday, March 12th, 2010 at 13:38 by David Derrick

Our link between Wolverhampton and Birmingham has gone down. We are investigating with our supplier. Traffic is being rerouted via alternate paths, some increase in latency may be seen.

  1. David Derrick says:

    Our supplier reports that engineers should be arriving shortly to find the exact location of the fibre break. There is no ETR at this time.

  2. Steve Lalonde says:

    Service was restored at 22:05 All traffic is now on the preferred paths.

At-Risk: Telford Data Centre 3 Cooling

Posted: Monday, March 8th, 2010 at 16:37 by Richard Partridge

We are currently aware of and working on an issue with one of our air cooling units in Data Centre 3 which appears to have failed. Measures are currently in place to minimise the impact that the loss of this unit is having.

Although we do not expect this to adversely affect the service we offer, this does mean that the remaining Data Centre cooling system is currently running at an increased load and we advise customers to consider the site “At Risk” until repair work has been completed.

Cooling engineers are working on the fault and the current ETR is 09/03/2010 12:00.

  1. Richard Partridge says:

    All repair work has now been completed and the cooling system is now operating at normal levels.